# Q-score Max-Clique: The First Quantum Metric Evaluation on Multiple Computational Paradigms

### v2

@@ 4.1 Experiments
    Figure 1: Max-Clique Q-score’s $\beta$ vs problem size $N$ for the D-wave Advantage and 2000Q QPUs. For visibility reasons, the advantage graph is off-set by $1$.
    Figure 2: Max-Clique Q-score’s $\beta$ vs problem size $N$, with respect to optimal clique solution $C_{max}$, for the D-wave Advantage and 2000Q QPUs. For visibility reasons, the advantage graph is off-set by $1$.
@@
  • Classical and hybrid: Lastly, Fig. 4 shows the $\beta$ vs $N$ graph for D-wave’s classical and hybrid solvers. For the classical solvers, the simulations start for $N=100$ and increase in size with steps of $100$. The computation is terminated when the average time for graph sizes exceeds $100$ seconds on our server to ensure that the region of 0 to $60$ seconds is captured completely. For the hybrid solver, $\beta$ is calculated for problem instances starting from $N=1{,}500$ up to $N=12{,}500$ with steps of $500$ and a time constraint of $60$ seconds. Due to the relative high cost of resources of the hybrid solver, only a single problem instance is considered per problem size.
@@
    Each of these solvers is able to maintain a $\beta>0.2$ for large problem instances $N$. Therefore, the Q-score Max-Clique is determined by the imposed 60 seconds time constraint. Fig. 5 shows the average time taken to solve a problem instance. For the hybrid solver, the minimum needed computation time is hard-coded. From these results, we report a Q-score Max-Clique of $9{,}100$ and $4{,}900$ for simulated annealing and tabu search respectively. The hybrid solver has a Q-score Max-Clique of $12{,}500$ because of the hard-coded computation time constraint.
@@
    Figure 4: Max-Clique Q-score’s $\beta$ vs problem size $N$ for the D-wave classical tabu search solver (qbsolv), Simulated Annealing and their hybrid solver.
@@
    Figure 5: Average time taken per problem instance $N$ for the D-wave classical tabu search solver (qbsolv), Simulated Annealing and the hybrid solver.
5 Discussion
In this work, it is demonstrated that the Q-score metric, as originally introduced by Atos, extends to other optimisation problems. To compute a corresponding Q-score, such an extension requires a definition of a random algorithm and a corresponding asymptotic approximation of the optimal solution. Our choice for Max-Clique is motivated by the existence of a clique search algorithm for photonic hardware. This allowed us to produce the first quantum metric evaluation comparing three different quantum paradigms, as well as classical solvers.
@@ 5 Discussion
Defining a Q-score metric inherently leaves a lot of freedom, even for a fixed optimisation problem. For instance, one could argue that our random algorithm choice is too naive resulting in a constant $C_{rand}$ value, which might seem strange for large problem instances. A different, yet naive, algorithm would be to start from a random node and grow it to a clique by iterating over all possible vertices. In the end, different random algorithms will result in different $\beta$ vs $N$ graphs and hence different Q-scores. However, it is expected that this does not significantly change the relative performance between devices, and hence does not change the relative Q-scores.
Other degrees of freedom are the considered time constraint, the allowed pre- and post-processing optimisation steps and the value $\beta^{*}=0.2$ and even the considered (optimisation) problem. On the one hand, this forms a weakness of the Q-score metric: users should be very specific on how the metric is defined, which might complicate a potential comparison between different works. On the other hand, this also allows for different ways of benchmarking devices using a Q-score metric. A next step for the Q-score would be to consider a Q-score in which these degrees of freedom can be changed dynamically. The Q-score could hence be turned into a plug-and-play metric or suite of metrics suitable for benchmarking devices in a wide variety of ways. Specifically, given a certain use case, the problem to define the Q-score can be chosen so that it is very relevant for the use case, hence allowing devices to be specifically benchmarked for the use case.
@@
In this work, the hybrid solver of D-Wave obtained the highest Q-score Max-Clique of $12{,}500$ of all considered solvers. The differences with other classical solvers are however not yet significant enough to hint towards a quantum advantage. The hybrid solver forms a black box, which hinders attributing the better performance solely to the quantum routines. Specifically, the hybrid solver uses a server stronger than our local solver for its classical routines. This inherently makes the comparison between the hybrid and classical solvers somewhat unfair. In addition, the $\beta$ vs $N$ graph for the hybrid solver shows a seemingly inexplicable drop at around $N=11000$. It seems that there is something in the black box workings of the hybrid solver that results in problems of size larger than $N=10000$ being relatively more difficult to solve for the hybrid solver.
@@
Comparing Q-score Max-Clique with earlier results for Max-Cut [14], shows that the Q-score for the Advantage system is slightly lower ($110$ versus $140$), while the 2000Q device finds the same score of 70 for both Q-scores. On the contrary, the classical approach performs better for the Max-Clique problem, with scores of $2{,}300$ vs $4{,}900$ and $5{,}800$ vs $9{,}100$ for tabu search and simulated annealing respectively. However, upon closer inspection that does not tell the whole story. Specifically, for the Q-score Max-Cut, the classical and hybrid solvers find solutions of perfect quality $\beta=1$ throughout, while for Q-score Max-Clique, the same solvers show both a variety and decrease in the solution quality parameter $\beta$. So, while the absolute Q-scores Max-Clique are higher, the Q-score Max-Cut finds higher quality solutions throughout. This hints towards the added value of a suite of Q-scores for different optimisation problems and parameter settings, for example using different $\beta^{*}$-values. In this way, the best hardware for each optimisation problem and application can be found.
For further work, it would be interesting to extend the QAOA results to more hardware backends and larger problem instances. For a fair comparison, additional research is needed on the execution time of QAOA. Alternatively, it would also be interesting to consider different gate-based algorithm approaches. For example, one could also solve the Max-Clique problem using a Grover’s search [39].

### v3

@@ IV-A Experiments
    Figure 1: Max-Clique Q-score’s $\beta$ vs problem size $N$ for the D-Wave Advantage and 2000Q QPUs. For visibility reasons, the Advantage graph is off-set by $1$ in the x-dimension.
    Figure 2: Max-Clique Q-score’s $\beta$ vs problem size $N$, with respect to optimal clique solution $C_{max}$, for the D-Wave Advantage and 2000Q QPUs. For visibility reasons, the Advantage graph is off-set by $1$ in the x-dimension.
@@
  • Classical and hybrid: Lastly, Fig. 4 shows the $\beta$ vs $N$ graph for D-Wave’s classical and hybrid solvers. For the classical solvers, the simulations start for $N=100$ and increase in size with steps of $100$. The computation is terminated when the average time for graph sizes exceeds $100$ seconds on our server to ensure that the region of 0 to $60$ seconds is captured completely. For the hybrid solver, $\beta$ is calculated for problem instances starting from $N=1{,}500$ up to $N=12{,}500$ with steps of $500$ and a time constraint of $60$ seconds. Due to the relative high cost of resources of the D-Wave hybrid solver, only a single problem instance is considered per problem size.
@@
    Each of these solvers is able to maintain a $\beta>0.2$ for large problem instances $N$. Therefore, the Q-score Max-Clique is determined by the imposed 60 seconds time constraint. Fig. 5 shows the average time taken to solve a problem instance. For the hybrid solver, the minimum needed computation time is hard-coded. From these results, we report a Q-score Max-Clique of $9{,}100$ and $4{,}900$ for the D-Wave implementations of simulated annealing and tabu search respectively. The hybrid solver has a Q-score Max-Clique of $12{,}500$ because of the hard-coded computation time constraint.
@@
    Figure 4: Max-Clique Q-score’s $\beta$ vs problem size $N$ for the D-Wave classical tabu search solver (qbsolv), D-Wave Simulated Annealing and their hybrid solver.
@@
    Figure 5: Average time taken per problem instance $N$ for the D-Wave classical tabu search solver (qbsolv), D-Wave Simulated Annealing and the hybrid solver.
V Conclusion and Discussion
In this work, the Q-score as originally introduced by Atos is extended to a framework definition. First, it is noted that the original Q-score definition and subsequent work make many arbitrary choices, which should actually be left as degrees of freedom for the user. By stripping the original Q-score of these arbitrary choices, the Q-score framework can be defined. In this framework, optimisation problem, data set, $\beta^{*}$-value, allowed runtime and allowed optimisation steps are left as degrees of freedom. By choosing these degrees of freedom to optimally match the use case at hand, a suitable metric can be designed for each use case. The Q-score framework then captures a score defined as the largest problem size for which a device on average significantly outperforms a random approach.
To showcase the applicability of the Q-score framework, this work also defines the Q-score Max-Clique. Our choice for Max-Clique is motivated by the existence of a clique search algorithm for photonic hardware. This allowed us to define the first quantum metric which can be natively implemented on three quantum paradigms, as well as classical computers. This metric is instantiated with specific degrees of freedom, after which it is applied to solvers from all four paradigms. This evaluation shows that it is indeed possible to use the Q-score Max-Clique to benchmark all paradigms.
It is important to note that there are other ways of defining quantum metrics that can compare all four paradigms. As an example, the BACQ initiative focuses on the original Q-score metric to benchmark photonic devices as well [46]. This can be done, as the Max-Cut problem can be rewritten as a problem that is solvable on photonic quantum devices, such as the Max-Clique problem. Both methods are valid ways of yielding a metric that is capable if benchmarking solvers of all paradigms, with neither being better than the other. One showcases the value of the Q-score as a framework, while the other showcases the strength of the original Q-score Max-Cut metric itself.
The strength of the Q-score framework comes from the fact that it contains many degrees of freedom that can be optimally matched to different use cases. It could be argued that this is also a weakness of the Q-score: users should be very specific on how the metric is defined, which might complicate a potential comparison between different works. Hardware vendors, in particular, have an incentive to select the degrees of freedom in a way that suit the performance of their hardware. This makes the Q-score framework less suitable to form "the" application-level metric that should be used in any situation. However, that is also not the goal of the framework: instead, it yields a method that has gives valuable insights for any use case.
@@ V Conclusion and Discussion
In the definition of the Q-score Max-Clique, various degrees of freedom were chosen in quite an arbitrary way. Of course, different choices could be made for the various degrees of freedom, some of which will be discussed here. Firstly, one could argue that our choice for random algorithm is too naive resulting in a constant $C_{rand}$ value, which might seem strange for large problem instances. A different, yet naive, algorithm would be to start from a random node and grow it to a clique by iterating over all possible vertices. In the end, different random algorithms will result in different $\beta$ vs $N$ graphs and hence different Q-scores. However, it is expected that this does not significantly change the relative performance between devices, and hence does not change the relative Q-scores.
Secondly, the runtime limit was chosen as 60 seconds. Although other choices would have been reasonable as well, this was decided following earlier work [18]. Unfortunately, this time limit was too high for the photonic and gate-based quantums solvers, mainly due to queueing times. To showcase results for these solvers as well, it was decided to perform the benchmark for these paradigms without the time constraint. It was decided to keep the time limit for the other solvers, as especially the classical solvers would have otherwise obtained unrealistically high Q-scores due to their brute-force abilities. Naturally, this makes the comparison using this benchmark somewhat unfair. However, as mentioned before, this is not an issue for this work, as the goal of this work is not to perform a fair comparison, but rather the showcase the applicability of the Q-score Max-Clique. Once the different paradigms have matured, and queueing times have decreased, it is expected that using the same time limit throughout will result in a fair benchmark.
@@
In this work, the hybrid solver of D-Wave obtained the highest Q-score Max-Clique of $12{,}500$ of all considered solvers. The differences with other classical solvers are however not yet significant enough to hint towards a quantum advantage. The hybrid solver forms a black box, which hinders attributing the better performance solely to the quantum routines. Unfortunately, D-Wave does not publish any information regarding these precise workings. What is known, however, is that the hybrid solver uses a server stronger than our local solver for its classical routines. Naturally, this yields an unfair comparison, and it is in particular also expected that the classical solvers will obtain higher scores if they were to be implemented using stronger servers. All in all, it can be concluded that the performed benchmark should not be used to compare the solvability of the hybrid solver to the classical solvers. Again, this is not an issue for this work. For future practical use of the Q-score Max-Clique, it is expected that both solvers will be benchmarked using similar server capacity. In addition, the $\beta$ vs $N$ graph for the hybrid solver shows a seemingly inexplicable drop at around $N=11000$. It seems that there is something in the black box workings of the hybrid solver that results in problems of size larger than $N=10000$ being relatively more difficult to solve for the hybrid solver.
@@
Comparing Q-score Max-Clique with earlier results for Max-Cut [18], shows that the Q-score for the Advantage system is slightly lower ($110$ versus $140$), while the 2000Q device finds the same score of 70 for both Q-scores. On the contrary, the classical approach performs better for the Max-Clique problem, with scores of $2{,}300$ vs $4{,}900$ and $5{,}800$ vs $9{,}100$ for the D-Wave implementations of tabu search and simulated annealing respectively. However, upon closer inspection that does not tell the whole story. Specifically, for the Q-score Max-Cut, the classical and hybrid solvers find solutions of perfect quality $\beta=1$ throughout, while for Q-score Max-Clique, the same solvers show both a variety and decrease in the solution quality parameter $\beta$. So, while the absolute Q-scores Max-Clique are higher, the Q-score Max-Cut finds higher quality solutions throughout. This hints towards the added value of a suite of Q-scores for different optimisation problems and parameter settings, for example using different $\beta^{*}$-values. In this way, the best hardware for each optimisation problem and application can be found.
For future work, it would naturally be most interesting to consider other instantiations of the Q-score framework. This should consider both other problems, as well as changing the other degrees of freedom. In addition, it would be interesting to consider how evaluations using the different Q-scores compare, and how these different Q-scores compare to other application-level metrics.
Regarding future work for the Q-score Max-Clique, it would be interesting to extend the QAOA results to more hardware backends and larger problem instances. For a fair comparison, additional research is needed on the execution time of QAOA. Alternatively, it would also be interesting to consider different gate-based algorithm approaches. For example, one could also solve the Max-Clique problem using a Grover’s search [47] or a variational quantum eigensolver [48].